# Pedestrian Tracking with Monocular Camera using Unconstrained 3D Motion Model

### v1

@@ 4.1.3 Physical Interpretation and Setup of Model Parameters
[…]
@@
4.2 Model-Based Filter for (4.1.2)
From (38a), note that the KF depends on the latest measurement linearly. If the state-space model is nonlinear and the KF cannot be applied, one can still define an optimal estimator with minimal MSE, subject to being linear w.r.t. the latest measurement, similarly to the KF. One way to approximately implement131313 Assuming the depth $z_{k}$ and its uncertainty is such that the planar 3D BB is sufficiently far from the singularity at $z_{k}=0$. the estimator is to use the unscented transform (UT) leading to the unscented Kalman filter (UKF) [25, 26].
4.2.1 Unscented Transform
Consider a random vector $\mathbf{x}$ of dimension $n_{\mathbf{x}}$ with mean $\mathbf{m}_{\mathbf{x}}$ and covariance matrix $\mathbf{P}_{\mathbf{x}}$, and a random vector $\mathbf{y}=\mathbf{f}(\mathbf{x})$ resulting as a nonlinear transformation of $\mathbf{x}$. The objective is to find an approximate mean $\mathbf{m}_{\mathbf{y}}$ and covariance matrix $\mathbf{P}_{\mathbf{y}}$ of $\mathbf{y}$. Instead of approximating the nonlinear function, the UT uses deterministically chosen sigma-points [25]. The sigma-points are computed as141414 For simplicity, only a symmetric sigma-point set without additional scaling is considered in this paper, but other formulations exist [25].
@@ 4.2.1 Unscented Transform
4.2.2 UKF Implementation of the Model-Based Filter
@@
The filtering step is given by the UKF update. Using a numerically stable covariance matrix update based on the square-root UKF implementation [26],
@@
If 2D BBs are needed in addition to the (predictive) estimates of the planar 3D BB states, one can directly output $\mathbf{m}_{\mathbf{y}_{k}}$ (85a) and $\mathbf{P}_{\mathbf{y}_{k}}=\mathbf{M}_{\mathbf{y}_{k}}\mathbf{M}_{\mathbf{y}_{k}}^{\top}$ (85b). Note that filtering 2D BB estimates are obtained analogically from $\mathbf{x}_{k|k}^{\mathrm{3D}}$ and $\mathbf{P}_{k|k}^{\mathrm{3D}}$.
4.2.3 Initialization Using a Measurement
@@
To take the measurement noise into account, the variables $\hat{\mathbf{r}}_{C,O}(\mathbf{z}_{0})=\mathbf{m}_{\mathbf{r}}$ (85a) and $\widehat{\mathbf{P}}_{C,O}(\mathbf{z}_{0})=\mathbf{P}_{\mathbf{r}}$ (85b) are computed using the UT with function $\mathbf{r}=\mathbf{f}(\mathbf{x};\mathbf{z}_{0})$, specified as follows. The random vector $\mathbf{x}$ is of dimension $n_{\mathbf{x}}=4$, has the mean $\mathbf{m}_{\mathbf{x}}=[\mathbf{0}_{1\times 3}\ m_{h}]^{\top}$ and covariance matrix $\mathbf{P}_{\mathbf{x}}=\text{diag}\left(\,{\mathbf{JRJ}^{\top},\sigma_{h}^{2}}\,\right)$, where $\mathbf{J}\!=\!\left[\begin{smallmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ \end{smallmatrix}\right]$ is designed to omit the measurement noise corresponding to the width. The function $\mathbf{f}(\,\cdot\,;\mathbf{z}_{0}):\mathbb{R}^{4}\rightarrow\mathbb{R}^{3}$ serves to compute an inverse of the noisy perspective projection for a 2D point $\left[[\mathbf{z}_{0}]_{1}\ [\mathbf{z}_{0}]_{2}\right]^{\top}$, i.e., of the first and third rows of (62b), by estimating the depth161616 From Section 3.1.4, recall that using the height is preferable to the width for estimating the depth due to possible hand swinging of a boxed pedestrian. similarly to (3.1.4), and it is given by
@@
5 Results With Real Data
@@
5.1 Retrieving of the 3D Information
As only annotations in 2D are available in the studied data subset, a reliable evaluation for entire state vectors is not possible. To arrive at indicative results, however, one can generate semi-annotations in 3D of the form of a planar 3D BBs by computing the depth similarly to (3.1.4), as $z_{k}\approx\tfrac{f\cdot h^{\mathrm{g}}}{\texttt{|px|}\cdot[\mathbf{a}_{k}]_{4}}$, where $h^{\mathrm{g}}$ in meters is the height of the given pedestrian guessed based on observing images. Applying (4)-(10), the semi-annotations are
@@ 5.2 Evaluation Metrics
5.3 Results for Real and Simulated Detections
@@
The legend of the following figures includes annotations or semi-annotations ([Uncaptioned image]); gray points ([Uncaptioned image]) indicating availability of detections; results with the 2D-Model-Based Filter (Section 3.1) ([Uncaptioned image]), Bag-of-Tricks-Based Filter (Section 3.2) ([Uncaptioned image]), and the proposed method ([Uncaptioned image]). The dotted lines ([Uncaptioned image]) indicate edges of the camera’s line of sight, and the stars ([Uncaptioned image]), ([Uncaptioned image]) and ([Uncaptioned image]) indicate medians computed over time-plots of the corresponding color.
5.3.1 Sequence MOT17-02, pedestrian with id 2
@@
[…]
Figure 4: Scenario 5.3.1: estimated individual state variables of the proposed filter with standard deviation error intervals (Refer to caption), semi-annotations (Refer to caption), and ANEES with $\mathbf{x}_{k}=[x_{k}\ y_{k}\ z_{k}\ \omega_{k}\ h^{\mathrm{g}}]^{\top}$ using Faster R-CNN detections.
Figure 5: Scenario 5.3.1: RMSE and ANEES with $\mathbf{x}_{k}=[\mathsf{x}_{k}\ \mathsf{y}_{k}\ \upomega_{k}\ \mathsf{h}_{k}]^{\top}$ and Faster R-CNN detections.
Figure 6: Scenario 5.3.1: RMSE and ANEES with $\mathbf{x}_{k}=[\mathsf{x}_{k}\ \mathsf{y}_{k}\ \upomega_{k}\ \mathsf{h}_{k}]^{\top}$ and simulated detections.
@@
5.3.2 Sequence MOT17-09, pedestrian with id 23
@@
Figure 7: Scenario 5.3.2: Results of the proposed filter with the F-RCNN detector, including error covariance ellipses (plotted every 20 ms).
Figure 8: Scenario 5.3.2: RMSE and ANEES with $\mathbf{x}_{k}=[\mathsf{x}_{k}\ \mathsf{y}_{k}\ \upomega_{k}\ \mathsf{h}_{k}]^{\top}$ and Faster R-CNN detections.
Figure 9: Scenario 5.3.2: RMSE and ANEES with $\mathbf{x}_{k}=[\mathsf{x}_{k}\ \mathsf{y}_{k}\ \upomega_{k}\ \mathsf{h}_{k}]^{\top}$ and simulated detections.
@@
The proposed method is approximately $3.6$ times slower than both of the baseline 2D filters. The Bag-of-Tricks-Based Filter is only about $15$% slower than the 2D-Model-Based Filter.
6 Conclusion

### v2

@@ IV-A3 Physical Interpretation and Setup of Model Parameters
which is approximately equal to 1 m2s-3 for $T=\nicefrac{{1}}{{30}}$ s and $\ddot{r}^{\mathrm{max}}=5.477$ $\cdots$-2, i.e., if the maximal change of the object’s speed over a second is $1.522$ km/h. The speed of pedestrians for the MOT-17 dataset are also analyzed in [19].
@@
IV-B Model-Based Filter for (33)
From (18a), note that the KF depends on the latest measurement linearly. If the state-space model is nonlinear and the KF cannot be applied, one can still define an optimal estimator with minimal MSE, subject to being linear w.r.t. the latest measurement, similarly to the KF. One way to approximately implement141414 Assuming the depth $z_{k}$ and its uncertainty is such that the planar 3D BB is sufficiently far from the singularity at $z_{k}=0$. the estimator is to use the unscented transform (UT) leading to the unscented Kalman filter (UKF) [25, 26].
IV-B1 Unscented Transform
Consider a random vector $\mathbf{x}$ of dimension $n_{\mathbf{x}}$ with mean $\mathbf{m}_{\mathbf{x}}$ and covariance matrix $\mathbf{P}_{\mathbf{x}}$, and a random vector $\mathbf{y}=\mathbf{f}(\mathbf{x})$ resulting as a nonlinear transformation of $\mathbf{x}$. The objective is to find an approximate mean $\mathbf{m}_{\mathbf{y}}$ and covariance matrix $\mathbf{P}_{\mathbf{y}}$ of $\mathbf{y}$. Instead of approximating the nonlinear function, the UT uses deterministically chosen sigma-points [25]. The sigma-points are computed as151515 For simplicity, only a symmetric sigma-point set without additional scaling is considered in this paper. As a result, the formulation herein is equivalent to the spherical cubature approximation, but other formulations exist [25].
@@ IV-B1 Unscented Transform
IV-B2 UKF Implementation of the Model-Based Filter
@@
The filtering step is given by the UKF update. Using a numerically stable covariance matrix update based on the square-root UKF implementation [26], it is
@@
If 2D BBs are needed in addition to the (predictive) estimates of the planar 3D BB states, one can directly output $\mathbf{m}_{\mathbf{y}_{k}}$ (40a) and $\mathbf{P}_{\mathbf{y}_{k}}=\mathbf{M}_{\mathbf{y}_{k}}\mathbf{M}_{\mathbf{y}_{k}}^{\top}$ (40b). Note that the filtering 2D BB estimates are obtained analogically from $\mathbf{x}_{k|k}^{\mathrm{3D}}$ and $\mathbf{P}_{k|k}^{\mathrm{3D}}$.
IV-B3 Initialization Using a Measurement
@@
To take the measurement noise into account, the variables $\hat{\mathbf{r}}_{C,O}(\mathbf{z}_{0})=\mathbf{m}_{\mathbf{r}}$ (40a) and $\widehat{\mathbf{P}}_{C,O}(\mathbf{z}_{0})=\mathbf{P}_{\mathbf{r}}$ (40b) are computed using the UT with function $\mathbf{r}=\mathbf{f}(\mathbf{x};\mathbf{z}_{0})$, specified as follows. The random vector $\mathbf{x}$ is of dimension $n_{\mathbf{x}}=4$, has the mean $\mathbf{m}_{\mathbf{x}}=[\mathbf{0}_{1\times 3}\ m_{h}]^{\top}$ and covariance matrix $\mathbf{P}_{\mathbf{x}}=\text{diag}\left(\,{\mathbf{JRJ}^{\top},\sigma_{h}^{2}}\,\right)$, where $\mathbf{J}\!=\!\left[\begin{smallmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ \end{smallmatrix}\right]$ is designed to omit the measurement noise corresponding to the width. The function $\mathbf{f}(\,\cdot\,;\mathbf{z}_{0}):\mathbb{R}^{4}\rightarrow\mathbb{R}^{3}$ serves to compute an inverse of the noisy perspective projection for a 2D point $\left[[\mathbf{z}_{0}]_{1}\ [\mathbf{z}_{0}]_{2}\right]^{\top}$, i.e., of the first and third rows of (33b), by estimating the depth171717 From Section III-A4, recall that using the height is preferable to the width for estimating the depth due to possible hand swinging of a boxed pedestrian. similarly to (21), and it is given by
@@
V Results With Real Data
@@
V-A Retrieving of the 3D Information
As only annotations in 2D are available in the studied data subset, a reliable evaluation for entire state vectors is not possible. To arrive at indicative results, however, one can generate semi-annotations in 3D of the form of a planar 3D BBs by computing the depth similarly to (21), as $z_{k}\approx\tfrac{f\cdot h^{\mathrm{g}}}{\texttt{|px|}\cdot[\mathbf{a}_{k}]_{4}}$, where $h^{\mathrm{g}}$ in meters is the height of the given pedestrian guessed based on observing images. Applying (2)-(6), the semi-annotations are
@@ V-B Evaluation Metrics
V-C Results for Real and Simulated Detections
@@
The legend of the following figures includes annotations or semi-annotations ([Uncaptioned image]); gray points ([Uncaptioned image]) indicating availability of detections; results with the 2D-Model-Based Filter (Section III-A) ([Uncaptioned image]), Bag-of-Tricks-Based Filter (Section III-B) ([Uncaptioned image]), and the proposed method ([Uncaptioned image]). The dotted lines ([Uncaptioned image]) indicate edges of the camera’s line of sight, and the stars ([Uncaptioned image]), ([Uncaptioned image]) and ([Uncaptioned image]) indicate medians computed over time-plots of the corresponding color.
V-C1 Sequence MOT17-02, pedestrian with id 2
@@
  • The proposed method ([Uncaptioned image]) provides consistent estimates as ANEESk fluctuates around one in both 2D and 3D. In contrast, the 2D-Model-Based Filter ([Uncaptioned image]) mostly provides optimistic estimates and the Bag-of-Tricks-Based Filter ([Uncaptioned image]) mostly provides pessimistic estimates.
Figure 4: Scenario V-C1: estimated individual state variables of the proposed filter with standard deviation error intervals (Refer to caption), semi-annotations (Refer to caption), and ANEES with $\mathbf{x}_{k}=[x_{k}\ y_{k}\ z_{k}\ \omega_{k}\ h^{\mathrm{g}}]^{\top}$ using Faster R-CNN detections.
Figure 5: Scenario V-C1: RMSE and ANEES with $\mathbf{x}_{k}=[\mathsf{x}_{k}\ \mathsf{y}_{k}\ \upomega_{k}\ \mathsf{h}_{k}]^{\top}$ and Faster R-CNN detections.
Figure 6: Scenario V-C1: RMSE and ANEES with $\mathbf{x}_{k}=[\mathsf{x}_{k}\ \mathsf{y}_{k}\ \upomega_{k}\ \mathsf{h}_{k}]^{\top}$ and simulated detections.
@@
V-C2 Sequence MOT17-09, pedestrian with id 23
@@
Figure 7: Scenario V-C2: Results of the proposed filter with the F-RCNN detector, including error covariance ellipses (plotted every 20 ms).
Figure 8: Scenario V-C2: RMSE and ANEES with $\mathbf{x}_{k}=[\mathsf{x}_{k}\ \mathsf{y}_{k}\ \upomega_{k}\ \mathsf{h}_{k}]^{\top}$ and Faster R-CNN detections.
Figure 9: Scenario V-C2: RMSE and ANEES with $\mathbf{x}_{k}=[\mathsf{x}_{k}\ \mathsf{y}_{k}\ \upomega_{k}\ \mathsf{h}_{k}]^{\top}$ and simulated detections.
@@
On average, the proposed method is about $10$ times slower than both baseline filters. The Bag-of-Tricks-Based Filter is only about $10$% slower than the 2D-Model-Based Filter.
VI Conclusion